# You Cannot Feed Two Birds with One Score: the Accuracy-Naturalness Tradeoff in Translation

### v1

@@ 1 Introduction
Machine translation (MT) is one of the undeniable success stories of computer science and machine learning research. Its history is long and varied, dating back to at least the 1960s and having progressed through symbolic, statistical and, most recently, neural approaches. By now, the performance of state-of-the-art systems is close to or matches that of human translators. But how do we assess the quality of translation systems? Indeed, how should we assess them? Already in the early days of MT, researchers recognized the need to evaluate translations along two axes: the ALPAC project considered the “intelligibility” and “fidelity” of the translations (Pierce & Carroll, 1966), while DARPA evaluations used “adequacy” and “fluency” (White & O’Connell, 1993). Although these concepts are intuitively appealing, they partly fell out of favour for two main reasons. First, they lack solid, widely-accepted formal definitions, which made it 1) difficult to define guidelines for human evaluations, and 2) impossible to establish their theoretical properties and their interaction. Second, in recent years, optimising solely some notion of accuracy where we compare an MT hypothesis with a given reference, has already lead to fluent translations in practice. Hence, even the relevance of such a distinction has become unclear. Nonetheless, this approach to automatic evaluation of machine translation has always had issues. For example, for a few years now, it has been clear that classical “overlap metrics,” such as BLEU and chrF are no longer good enough to assess translation quality, as they do not correlate well with human ratings, which prompted researchers to move towards neural metrics instead (Freitag et al., 2022a). However, in the latest WMT general task, a similar phenomenon occurred: systems with the best automatic scores (based on neural metrics) did not achieve the best score among human raters (Kocmi et al., 2024). This and related phenomena motivated us to reexamine translation evaluation practices.
@@
As we note above, at present there is an implicitly held belief (or wishful thought) in the community that there should exist some metric, which can be optimized on to yield a translation system that is simultaneously accurate and natural-sounding. To examine the plausibility of this belief, we propose formal definitions of accuracy and naturalness by extending the mature, and widely-accepted information-theoretic framework of Blau & Michaeli (2018). Then, we show that under these definitions, the above goal is hopeless: there is no metric, optimizing which will yield a system that is simultaneously accurate and perfectly natural-sounding. Our contributions are:
@@ 2 Background and Notation
The primary measure of a translator’s quality is their accuracy, that is, how well their translations capture the meaning of the source text. Therefore, quantifying accuracy necessarily involves a comparison between the translator’s output and a reference. In practice, the reference could be the source text, some translation produced by a professional translator that is taken to be ground truth, or both. In this paper, we refer to any metric that performs such a comparison as a reference metric. Classic examples include BLEU (Papineni et al., 2002) and chrF (Popović, 2015), which compare candidates with reference translations. They are convenient and simple, because they assess translations on a lexical level with the hope that proximity to the reference translation will ensure that the translations are correct at the semantic level as well. In recent years, the community has moved toward neural reference metrics, such as MetricX (Juraska et al., 2024) and Comet (Rei et al., 2020), which aim to assess translations directly at the semantic level instead. These metrics typically aim to directly predict human opinion scores from large language model (LLM) features. They still rely on reference translations, but are more robust and hence better suited for evaluating modern MT systems (Freitag et al., 2022a). As a further development, quality estimation (QE) metrics drop the need for a reference translation.
@@
Producing fluent translations has always been an objective of translation systems. In the case of MT, early system struggled to even produce grammatically correct sentences. As the quality of automatic systems improved, this became less of an issue, but the community encountered another problem, long known in the (human) translation community: translated text often does not sound completely natural, it contains translationese (Gellerstam, 1986). To study and eliminate this phenomenon, researchers have proposed several candidate metrics for naturalness, such as the language model marginal (log-)probability (Freitag et al., 2022b; Lim et al., 2024) as well as linguistically motivated metrics as proposed by Vanmassenhove et al. (2021). These scores are also called no-reference metrics as they do not rely on additional data. Note that no-reference metrics are not to be confused with QE metrics, which we consider to be reference metrics as we explained above. However, although they are intuitive, no-reference metrics lack a more formal mathematical justification. In this paper, we develop an abstract notion of naturalness based on information theory, and show how no-reference metrics conform to our definitions, which provides a unifying framework.
@@
We denote a randomly selected source text by the random variable ${\mathbf{x}}\in\mathcal{X}$ and a random target text by ${\mathbf{y}}\in\mathcal{Y}$, with $\mathcal{X}$ ($\mathcal{Y}$) the set of all possible source (target) texts. We denote probability distributions by the capital roman letters $P,Q$ and $R$, and denote expectation of a function $f$ with respect to distribution $P$ as $\mathbb{E}_{{\mathbf{x}}\sim P}[f({\mathbf{x}})]$. Source-target pairs are drawn from the joint distribution $P_{{\mathbf{x}},{\mathbf{y}}}$. This joint distribution also immediately defines the marginal over the source sentences $P_{\mathbf{x}}$ and marginal over the target sentences $P_{\mathbf{y}}$. For the purposes of this paper, we think of $P_{{\mathbf{x}}}$ as a “true” monolingual distribution over the source language and $P_{{\mathbf{y}}\mid{\mathbf{x}}}$ as the distribution of human translations into the target language given the source sentence ${\mathbf{x}}$. Therefore, since we obtain $P_{\mathbf{y}}$ by marginalising over the input, it is the distribution over “human translations into the target language.” Thus, it will be useful at times to consider a different distribution $R_{\mathbf{y}}$, which we can think of as a “true” monolingual reference over the target language. Finally, we model a translation system as a conditional distribution $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$. Then, we can also immediately consider the “translator’s marginal” $Q_{\mathbf{y}}(\cdot)=\mathbb{E}_{{\mathbf{x}}\sim P_{\mathbf{x}}}[Q_{{\mathbf{y}}\mid{\mathbf{x}}}(\cdot)]$.
@@ 3 The Accuracy-Naturalness Tradeoff in Translation
We start by considering a direct transfer of the distortion metrics defined in Blau & Michaeli (2018). These are functions ${\Delta:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}}$ that take as input a reference translation $y^{r}$ and a candidate translation $y^{c}$ and output a real number that measures how close the candidate translation is to the reference translation. We require that: 1) $\Delta$ be bounded from below and 2) $\Delta$ is minimised when $y^{c}$ equals $y^{r}$. Lexical metrics like (negative) BLEU and (negative) ChrF clearly fulfill these conditions, but we need to expand this definition to accommodate modern MT evaluation metrics. First we note that some neural metrics (including QE metrics) take the source sentence as an additional “reference”. We thus expand our functional form to ${\Delta:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}}$. The interpretation of the output value should also be generalized to the degree in which the candidate $y^{c}$ captures the meaning with respect to $x$ and $y^{r}$, and thus it is minimized when $y^{c}$ is semantically equivalent to the given reference. As such, a reference metric corresponds to a notion of inaccuracy. Note that condition 2) is difficult to prove for neural metrics, but it certainly is in the spirit of their design. In most cases the reference metrics are “accuracy-like” (higher is better), which can be considered just as a negated distortion, and thus we define the accuracy (or reference-score) of a translation system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$ as:
$A(Q_{{\mathbf{y}}\mid{\mathbf{x}}})=-\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}})=-\mathbb{E}_{{\mathbf{x}},{\mathbf{y}}^{r}\sim P_{{\mathbf{x}},{\mathbf{y}}}}[\mathbb{E}_{{\mathbf{y}}^{c}\sim Q_{{\mathbf{y}}\mid{\mathbf{x}}}}[\Delta({\mathbf{x}},{\mathbf{y}}^{r},{\mathbf{y}}^{c})]]$ (1)
Virtually all commonly used translation metrics are designed to be reference metrics, they differ mainly on what arguments they utilize. For example, BLEU and ChrF ignore the source, QE metrics ignore the target reference, and MetricX  [^1] In their latest version.or Comet use all three.
@@
Following Blau & Michaeli (2018), we argue that the naturalness of a translated text should be intrinsic to it, and should not depend on the source text. The basis of our thesis is the following desideratum: a set of translations are perfectly natural if no observer/critic can reliably distinguish them from a monolingual corpus in the target language. More precisely, let $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$ denote a (possibly and usually stochastic) translator. Then, we consider the “translator’s marginal” $Q_{\mathbf{y}}(\cdot)=\mathbb{E}_{{\mathbf{x}}\sim P_{\mathbf{x}}}[Q_{{\mathbf{y}}\mid{\mathbf{x}}}(\cdot)]$, which is the output distribution we obtain when we randomly select an input ${\mathbf{x}}\sim P_{\mathbf{x}}$ and translate it: ${\mathbf{y}}\sim Q_{{\mathbf{y}}\mid{\mathbf{x}}}$. Then, we say that a translation system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$ is perfectly natural with respect to a monolingual reference distribution $R_{\mathbf{y}}$ in the target language, if the system’s marginal matches the reference: $Q_{\mathbf{y}}=R_{\mathbf{y}}$. Crucially, it is up to the practitioner to define $R_{\mathbf{y}}$. Of course, one could set $R_{\mathbf{y}}=P_{\mathbf{y}}$ to the marginal defined by our parallel dataset; however, this is not necessary and indeed there are many good reasons not to! For example, the reference translations might contain “human translationese” (Gellerstam, 1986) and it is usually better to assess the quality of the system against text that was conceived in the target language. Furthermore, translation system for different domains all require different notions of “naturalness” due e.g. to the specific terminology that they use. We might also wish to select $R_{\mathbf{y}}$ such that its support is filtered for unwanted content, such as slurs/curse words.
Requiring perfectly natural translations is in practice too stringent and hence we relax it such that $Q_{\mathbf{y}}\approx R_{\mathbf{y}}$ instead, which we formalise using statistical distances $D(Q_{\mathbf{y}},R_{\mathbf{y}})$. We require that $D$ be a divergence, in the sense that $D(Q,P)\geq 0$ for any two distributions $Q,P$ and $D(Q,P)=0$ if and only if $P=Q$. However, similarly to distortion, a statistical distance measures the unnaturalness or disfluency of a system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$. Hence, we define the naturalness with respect to a statistical distance $D$ and reference distribution $R_{\mathbf{y}}$ as the negated distance
@@
What distance $D$ should we choose? Natural choices of $D$ include integral probability metrics (IPM) such as the total variation and Wasserstein distances, and $f$-divergences, such as the Kullback-Leibler or Rényi divergences, as these families all have natural connections to distinguishability (Sriperumbudur et al., 2009), see Appendix A for a discussion. Finally, note that our proposed theory has no notion of “sentence-level” naturalness. Indeed, there are many examples where a translation is perfectly accurate and also sounds fluent. We instead focus on the naturalness as a property of the full system’s output distribution.
@@ 3.1 The Tradeoff
However, if we have a system that is perfectly accurate, can we ever expect it to be perfectly natural? To answer this, consider neural metrics, such as MetricX or Comet, which are trained so that they implicitly jointly assess adequacy and fluency (in the MT sense). Concretely, let $\Delta^{*}$ be a neural metric we obtain by optimizing some appropriate objective. Now, consider a translation system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ that minimizes $\Delta^{*}$, i.e., ${\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*})=\min_{Q_{{\mathbf{y}}\mid{\mathbf{x}}}}\Delta^{*}(Q_{{\mathbf{y}}\mid{\mathbf{x}}})}$. By definition, $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ is perfectly accurate, but when is it also perfectly natural? Letting ${Q_{\mathbf{y}}^{*}(\cdot)=\mathbb{E}_{{\mathbf{x}}\sim P_{\mathbf{x}}}[Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}(\cdot)}]$, we see that by our definition, $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}(\cdot)$ is perfectly natural when $D(Q_{\mathbf{y}}^{*},R_{\mathbf{y}})=0$, which occurs if and only if $Q_{\mathbf{y}}^{*}=R_{\mathbf{y}}$! Due to the flexibility of our framework, this allows us to construct a perfectly natural system from a perfectly accurate one: we just need to set the monolingual reference distribution $R_{\mathbf{y}}\leftarrow Q_{\mathbf{y}}^{*}$! However, this choice makes $R_{\mathbf{y}}$, and hence our notion of naturalness depend on $P_{{\mathbf{x}},{\mathbf{y}}}$, which we argue is problematic for assessing translation naturalness. Our philosophy is that the notion of naturalness in a given language $\mathcal{Y}$ should not depend implicitly or explicitly on another language $\mathcal{X}$. Otherwise, we arrive at the absurd conclusion that the naturalness in language $\mathcal{Y}$ is different when translating from language $\mathcal{X}_{1}$ into $\mathcal{Y}$ and when translating from language $\mathcal{X}_{2}$ into $\mathcal{Y}$. However, $Q_{\mathbf{y}}^{*}$ breaks our principle in two ways: 1) it depends on $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$, which in turn depends on $P_{{\mathbf{x}},{\mathbf{y}}}$ and 2) it depends the marginalisation. On the other hand, setting the reference distribution $R_{\mathbf{y}}$ to anything other than $Q_{\mathbf{y}}^{*}$ derived from a minimizer of $\Delta^{*}$ means that $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ will not be perfectly natural. This argument illuminates that if we follow the philosophy that the monolingual reference $R_{\mathbf{y}}$ should not depend on $P_{{\mathbf{x}},{\mathbf{y}}}$, then the degenerate case of perfect accuracy implying perfect naturalness is exceedingly unlikely in practice. However, we emphasize that this conclusion relies on explictly disallowing $R_{\mathbf{y}}$ to depend on $P_{{\mathbf{x}},{\mathbf{y}}}$ the motivation for which comes from the theory’s application to translation, and in some other scenarios allowing such dependence might be sensible. Note that similar results might hold in different settings too: in Appendix B we show a generalisation of Theorem 1 of Blau & Michaeli (2018), which does allow $P_{\mathbf{y}}$-dependence and admits a similar result.
@@
The above argument guarantees that there is almost always a tradeoff between accuracy and naturalness, but gives no estimate of its severity. Indeed, we might expect the tradeoff to be mild for language pairs with an almost one-to-one correspondence, such as Spanish-Catalan or Serbian-Croatian. On the other hand, the tradeoff should be more pronounced for more distant language pairs such as German-Japanese or Hungarian-Swahili. We quantify these tradeoffs in Section 5, specifically in Figure 3. Hence, to study the tradeoff, we define the accuracy-naturalness (AN) function [^2] The accuracy-naturalness function is related to the better-known distortion-perception function ${\delta(P)=\min_{Q_{{\mathbf{y}}\mid{\mathbf{x}}}}\{\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}})\}\text{ s.t. }D(Q_{\mathbf{y}},P_{\mathbf{y}})\leq P}$ (Blau & Michaeli, 2018) via $\delta(P)=-A(-P)$.
$A(N)=\max_{Q_{{\mathbf{y}}\mid{\mathbf{x}}}}\{-\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}})\}\quad\text{subject to }-D(Q_{\mathbf{y}},P_{\mathbf{y}})\geq N.$ (3)
@@
$A(N)$ represents the best accuracy we can ever expect for a translation system, subject to achieving a naturalness of at least $N$. We have the following result, analogous to Theorem 2 of Blau & Michaeli (2018); see Appendix C for the proof.
@@
Let $A(N)$ be a distortion-perception function defined by distortion $\Delta$, distributional distance $D$, parallel distribution $P_{{\mathbf{y}},{\mathbf{x}}}$ and monolingual reference $R_{\mathbf{y}}$. Then:
@@
That $A(N)$ is non-increasing in $N$, taken together with our argument above that a tradeoff exists for virtually all practically relevant cases implies that close to the curve, accuracy and naturalness must anti-correlate. I.e. if we want to increase the naturalness of a system, its accuracy must forcibly drop (and the other way round). The convexity of $A(N)$ in $N$ on the other hand means that we can expect a larger and larger degradation in naturalness for a small gain in accuracy and vice versa, especially at the extremes.

### v3

@@ 1 Introduction
Machine translation (MT) is one of the undeniable success stories of computer science and machine learning research. Its history is long and varied, dating back to at least the 1960s and having progressed through symbolic, statistical and, most recently, neural approaches. By now, the performance of state-of-the-art systems is close to or matches that of human translators. But how do we assess the quality of translation systems? Indeed, how should we assess them? Already in the early days of MT, researchers recognized the need to evaluate translations along two axes: the ALPAC project considered the “intelligibility” and “fidelity” of the translations (Pierce & Carroll, 1966), while DARPA evaluations used “adequacy” and “fluency” (White & O’Connell, 1993). Although these concepts are intuitively appealing, they partly fell out of favour for two main reasons. First, they lack solid, widely-accepted formal definitions, which made it 1) difficult to define guidelines for human evaluations, and 2) impossible to establish their theoretical properties and their interaction. Second, in recent years, optimising solely some notion of accuracy where we compare an MT hypothesis with a given reference, has already lead to fluent translations in practice. Hence, even the relevance of such a distinction has become unclear. Nonetheless, using a single metric for automatic translation evaluation has always had issues. For example, for a few years now, it has been clear that classical “overlap metrics,” such as Bleu and chrF are no longer good enough to assess translation quality, as they do not correlate well with human ratings, which prompted researchers to move towards neural metrics instead (Freitag et al., 2022a). However, in the latest WMT general task, a similar phenomenon occurred: systems with the best automatic scores (based on neural metrics) did not achieve the best score among human raters (Kocmi et al., 2024). This and related phenomena motivated us to reexamine translation evaluation practices.
@@
As we note above, at present there is an implicitly held belief (or wishful thought) in the community that there should exist some metric, which can be optimized on to yield a translation system that is simultaneously accurate and natural-sounding. To examine the plausibility of this belief, we propose formal definitions of accuracy and naturalness by extending the mature, and widely-accepted information-theoretic framework of Blau & Michaeli (2018). Then, we show that under these definitions, the above goal is hopeless: there is no metric, optimizing which will yield a system that is simultaneously accurate and perfectly natural-sounding. Specifically, our contributions are:
@@ 2 Background and Notation
The primary measure of a translator’s ability/quality is their accuracy, that is, how well their translations capture the meaning of the source text. Therefore, quantifying accuracy necessarily involves a comparison between the translator’s output and a reference. In practice, the reference could be the source text, some translation produced by a professional translator that is taken to be ground truth, or both. In this paper, we refer to any metric that performs such a comparison as a reference metric. Classic examples include Bleu (Papineni et al., 2002) and chrF (Popović, 2015), which compare candidates with reference translations. They are convenient and simple, because they assess translations on a lexical level with the hope that proximity to the reference translation will ensure that the translations are correct at the semantic level as well. In recent years, the community has moved toward neural reference metrics, such as MetricX (Juraska et al., 2024) and Comet (Rei et al., 2020), which aim to assess translations directly at the semantic level instead. These metrics typically aim to directly predict human opinion scores from large language model (LLM) features. They still rely on reference translations, but are more robust and hence better suited for evaluating modern MT systems (Freitag et al., 2022a). As a further development, quality estimation (QE) metrics drop the need for a reference translation, but as they take the source text as input, we consider them reference metrics too.
@@
Producing fluent translations has always been an objective of translation systems. In the case of MT, early systems struggled to even produce grammatically correct sentences. As the quality of automatic systems improved, this became less of an issue, but the community encountered another problem, long known in the (human) translation community: translated text often does not sound completely natural, it contains translationese (Gellerstam, 1986). To study and eliminate this phenomenon, researchers have proposed several candidate metrics for naturalness, such as the language model marginal (log-)probability (Freitag et al., 2022b; Lim et al., 2024) as well as linguistically motivated metrics as proposed by Vanmassenhove et al. (2021). These scores are also called no-reference metrics as they do not rely on additional data. Note that no-reference metrics are not to be confused with QE metrics, which we consider to be reference metrics as we explained above. However, although they are intuitive, no-reference metrics lack a more formal mathematical justification. In this paper, we develop an abstract notion of naturalness based on information theory, and show how no-reference metrics conform to our definitions, which provides a unifying framework.
@@
We denote some randomly selected source text by the random variable ${\mathbf{x}}\in\mathcal{X}$ and some random target text by ${\mathbf{y}}\in\mathcal{Y}$, with $\mathcal{X}$ ($\mathcal{Y}$) the set of all possible source (target) texts. We denote probability distributions by the capital roman letters $P,Q$ and $R$, and denote the expectation of a function $f$ with respect to distribution $P$ as $\mathbb{E}_{{\mathbf{x}}\sim P}[f({\mathbf{x}})]$. Source-target pairs are drawn from the joint distribution $P_{{\mathbf{x}},{\mathbf{y}}}$. This joint distribution also immediately defines the marginal over the source sentences $P_{\mathbf{x}}$ and marginal over the target sentences $P_{\mathbf{y}}$. For the purposes of this paper, we think of $P_{{\mathbf{x}}}$ as a “true” monolingual distribution over the source language and $P_{{\mathbf{y}}\mid{\mathbf{x}}}$ as the distribution of human translations into the target language given the source sentence $\mathbf{x}$}. This philosophy reflects current practices in machine translation: the source language is typically derived from a monolingual corpus, which professional translators then translate into the target language. Indeed, there has been a recent explicit effort in the MT community to avoid back-translations or translations in both directions (Toral et al., 2018; Läubli et al., 2020).
Since we obtain $P_{\mathbf{y}}$ by marginalising over the input, it is the distribution over “human translations into the target language.” Thus, it will be useful at times to consider a different distribution $R_{\mathbf{y}}$, which we can think of as a “true” monolingual reference over the target language. Finally, we model a translation system as a conditional distribution $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$. Then, we can also immediately consider the “translator’s marginal” $Q_{\mathbf{y}}(\cdot)=\mathbb{E}_{{\mathbf{x}}\sim P_{\mathbf{x}}}[Q_{{\mathbf{y}}\mid{\mathbf{x}}}(\cdot)]$.
@@ 3 The Accuracy-Naturalness Tradeoff in Translation
We start by considering a direct transfer of the distortion metrics defined in Blau & Michaeli (2018). These are functions $\Delta:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$} that take as input a reference translation $y^{r}$ and a candidate translation $y^{c}$ and output a real number that measures how close the candidate translation is to the reference translation. We require that: 1) $\Delta$ be bounded from below and 2) $\Delta$ is minimised when $y^{c}$ equals $y^{r}$. Lexical metrics like (negative) Bleu and (negative) ChrF clearly fulfill these conditions, but we need to expand this definition to accommodate modern MT evaluation metrics. First we note that some neural metrics (including QE metrics) take the source sentence as an additional “reference”. We thus expand our functional form to $\Delta:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$}. The interpretation of the output value should also be generalized to the degree in which the candidate $y^{c}$ captures the meaning with respect to $x$ and $y^{r}$, and thus it is minimized when $y^{c}$ is semantically equivalent to the given reference. As such, a reference metric corresponds to a notion of inaccuracy. Current neural metrics are known or at least assumed to violate this requirement. For example, they might be vulnerable to “universal translations,” which yield low distortion regardless of the reference (Yan et al., 2023). However, as we will see, our theory only requires the requirement to hold “on average,” which is indeed the case for neural metrics. In most cases, the reference metrics the community uses are “accuracy-like” (higher is better), which can be considered just as a negated distortion, and thus we define the accuracy (or reference-score) of a translation system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$ as:
$A(Q_{{\mathbf{y}}\mid{\mathbf{x}}})=-\mathbb{E}_{{\mathbf{x}},{\mathbf{y}}^{r}\sim P_{{\mathbf{x}},{\mathbf{y}}}}[\mathbb{E}_{{\mathbf{y}}^{c}\sim Q_{{\mathbf{y}}\mid{\mathbf{x}}}}[\Delta({\mathbf{x}},{\mathbf{y}}^{r},{\mathbf{y}}^{c})]]$ (1)
Virtually all commonly used translation metrics are designed to be reference metrics, they differ mainly on what arguments they utilize. For example, Bleu and ChrF ignore the source, QE metrics ignore the target reference, and MetricX  [^1] In their latest version.or Comet use all three. We note that in the community there is an unwritten feeling that neural metrics place a heavier weight on how well a particular translated sentence sounds than on the accuracy of the translation. However, so long as there is an element of comparison with a reference, and assuming that our assumptions are satisfied, neural metrics are still reference metrics.
@@
Following Blau & Michaeli (2018), we argue that the naturalness of a translated text should be intrinsic to it, and should not depend on the source text. The basis of our thesis is the following desideratum: a set of translations ought to be perfectly natural if no observer/critic can reliably distinguish them from a monolingual corpus in the target language. More precisely, let $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$ denote a (possibly and usually stochastic) translator. Then, we consider the “translator’s marginal” $Q_{\mathbf{y}}(\cdot)=\mathbb{E}_{{\mathbf{x}}\sim P_{\mathbf{x}}}[Q_{{\mathbf{y}}\mid{\mathbf{x}}}(\cdot)]$, which is the output distribution we obtain when we randomly select an input ${\mathbf{x}}\sim P_{\mathbf{x}}$ and translate it: ${\mathbf{y}}\sim Q_{{\mathbf{y}}\mid{\mathbf{x}}}$. Then, we say that a translation system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$ is perfectly natural with respect to a monolingual reference distribution $R_{\mathbf{y}}$ in the target language, if the system’s marginal matches the reference: $Q_{\mathbf{y}}=R_{\mathbf{y}}$. Crucially, it is up to the practitioner to define $R_{\mathbf{y}}$. Of course, one could set $R_{\mathbf{y}}=P_{\mathbf{y}}$ to the marginal defined by our parallel dataset; however, this is not necessary and indeed there are many good reasons not to! For example, the reference translations might contain “human translationese” (Gellerstam, 1986) and it is usually better to assess the quality of the system against text that was conceived in the target language. Furthermore, translation systems for different domains all require different notions of “naturalness” due, e.g., to the specific terminology that they use. We might also wish to select $R_{\mathbf{y}}$ such that its support is filtered for unwanted content, such as slurs/curse words.
In practice, requiring perfectly natural translations is too stringent and hence we relax it such that $Q_{\mathbf{y}}\approx R_{\mathbf{y}}$ instead, which we formalise using statistical distances. Concretely, we pick some divergence $D(Q,R)$, that is, $D(Q,R)\geq 0$ for any two distributions $Q,R$ and $D(Q,R)=0$ if and only if $Q=R$. However, similarly to distortion, a statistical distance measures the unnaturalness or disfluency of a system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$. Hence, we define the naturalness with respect to a statistical distance $D$ and reference distribution $R_{\mathbf{y}}$ as the negated distance
@@
What distance $D$ should we choose? Natural choices of $D$ include integral probability metrics (IPM) such as the total variation and Wasserstein distances, and $f$-divergences, such as the Kullback-Leibler or Rényi divergences, as these families all have natural connections to distinguishability (Sriperumbudur et al., 2009); see Appendix˜A for a discussion.
Finally, we emphasize that a key feature of our theory is to recognize naturalness as a “corpus-/dataset-level” property. As such, the theory has no notion of “sentence-level” naturalness. In other words, our theory concerns itself with the naturalness of a translation system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}$, but not with the naturalness of individual text segments.
@@ 3.1 The Tradeoff
However, if we have a system that is optimally accurate, can we ever expect it to be perfectly natural? To answer this, consider neural metrics, such as MetricX or Comet, which are trained so that they implicitly jointly assess adequacy and fluency (in the MT sense). Concretely, let $\Delta^{*}$ be a neural metric we obtain by optimizing some appropriate objective. Now, consider a translation system $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ that minimizes $\Delta^{*}$, i.e., $\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*})=\min_{Q_{{\mathbf{y}}\mid{\mathbf{x}}}}\Delta^{*}(Q_{{\mathbf{y}}\mid{\mathbf{x}}})$}. By definition, $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ is optimally accurate, but when is it also perfectly natural? Letting $Q_{\mathbf{y}}^{*}(\cdot)=\mathbb{E}_{{\mathbf{x}}\sim P_{\mathbf{x}}}[Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}(\cdot)$}], we see that by our definition, $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}(\cdot)$ is perfectly natural when $D(Q_{\mathbf{y}}^{*},R_{\mathbf{y}})=0$, which occurs if and only if $Q_{\mathbf{y}}^{*}=R_{\mathbf{y}}$! Due to the flexibility of our framework, this allows us to construct a perfectly natural system from a perfectly accurate one: we just need to set the monolingual reference distribution $R_{\mathbf{y}}\leftarrow Q_{\mathbf{y}}^{*}$! However, this choice makes $R_{\mathbf{y}}$, and hence our notion of naturalness depend on $P_{{\mathbf{x}},{\mathbf{y}}}$, which we argue is problematic for assessing translation naturalness. Our philosophy is that the notion of naturalness in a given language $\mathcal{Y}$ should not depend implicitly or explicitly on another language $\mathcal{X}$. Otherwise, we arrive at the absurd conclusion that the naturalness in language $\mathcal{Y}$ is different when translating from language $\mathcal{X}_{1}$ into $\mathcal{Y}$ and when translating from language $\mathcal{X}_{2}$ into $\mathcal{Y}$! However, $Q_{\mathbf{y}}^{*}$ breaks our principle in two ways: 1) it depends on $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$, which in turn depends on $P_{{\mathbf{x}},{\mathbf{y}}}$ and 2) it depends the marginalisation. On the other hand, setting the reference distribution $R_{\mathbf{y}}$ to anything other than $Q_{\mathbf{y}}^{*}$ derived from a minimizer of $\Delta^{*}$ means that $Q_{{\mathbf{y}}\mid{\mathbf{x}}}^{*}$ will not be perfectly natural. This argument illuminates that if we follow the philosophy that the monolingual reference $R_{\mathbf{y}}$ should not depend on $P_{{\mathbf{x}},{\mathbf{y}}}$, then the degenerate case of perfect accuracy implying perfect naturalness is exceedingly unlikely in practice. However, we emphasize that this conclusion relies on explictly disallowing $R_{\mathbf{y}}$ to depend on $P_{{\mathbf{x}},{\mathbf{y}}}$, the motivation for which comes from the theory’s application to translation. In some other scenarios allowing such dependence might be sensible: for example, in Appendix˜B, we generalise Theorem 1 of Blau & Michaeli (2018), which does allow $P_{\mathbf{y}}$-dependence and admits a similar “no-two-birds-with-one-scone” result.
@@
The above argument guarantees that there is almost always a tradeoff between accuracy and naturalness, but gives no estimate of its severity. Indeed, we might expect the tradeoff to be mild for language pairs with an almost one-to-one correspondence, such as Spanish-Catalan or Serbian-Croatian. On the other hand, the tradeoff should be more pronounced for more distant language pairs such as German-Japanese or Hungarian-Swahili. We quantify these tradeoffs in Section˜5, specifically in Figure˜3. Hence, to study the tradeoff, we define the accuracy-naturalness (AN) function [^2] The accuracy-naturalness function is related to the better-known distortion-perception function $\delta(P)=\min_{Q_{{\mathbf{y}}\mid{\mathbf{x}}}}\{\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}})\}\text{ s.t. }D(Q_{\mathbf{y}},R_{\mathbf{y}})\leq P$} (Blau & Michaeli, 2018) via $\delta(P)=-A(-P)$.
$A(N)=\max_{Q_{{\mathbf{y}}\mid{\mathbf{x}}}}\{-\Delta(Q_{{\mathbf{y}}\mid{\mathbf{x}}})\}\quad\text{subject to }-D(Q_{\mathbf{y}},R_{\mathbf{y}})\geq N.$ (3)
@@
$A(N)$ represents the best accuracy we can ever expect for a translation system, subject to achieving a naturalness of at least $N$. We have the following result, analogous to Theorem 2 of Blau & Michaeli (2018); see Appendix˜C for the proof.
@@
Let $A(N)$ be the accuracy-naturalness function defined by distortion $\Delta$, distributional distance $D$, parallel distribution $P_{{\mathbf{x}},{\mathbf{y}}}$ and monolingual reference $R_{\mathbf{y}}$. Then:
@@
That $A(N)$ is non-increasing in $N$, taken together with our argument above that a tradeoff exists for virtually all practically relevant cases implies that close to the curve, accuracy and naturalness must anti-correlate. That is, an increase in the naturalness of a system beyond a certain point must come at the sacrifice of some accuracy, and vice versa. The concavity of $A(N)$ in $N$, on the other hand, means that we can expect a larger and larger degradation in naturalness for a small gain in accuracy and vice versa, especially at the extremes.